# 3D Gaussian Parametric Head Model

### v1

@@ 4.1 Datasets
NeRSemble dataset contains over 260 different identities, and collects 72fps multi-view videos from 16 synchronized cameras for each identity. The total length of the videos of a single identity is approximately 6000-11000 frames. In the experiment, we selected 140 of the identities for training and the rest for evaluation. For each identity video, we selected about 150 frames from all 16 views as training data.
NPHM dataset contains 5200 3D human head scans. These scans come from 255 different identities, each with about 20 different expressions. We selected approximately 1600 scans of 80 identities for training. Since our method utilizes 2D images as training supervision, we render each scan from 80 different views to generate synthetic image data and record the camera parameters and the masks.
FaceVerse dataset is an East Asian human head scan dataset. It contains 2310 scans from 110 different identities, and each identity contains 21 expressions. We selected 1620 scans data of 80 identities for training. Similarly, for each scan, we render multi-view synthetic image data from 80 different views and record the camera parameters and the masks.
4.2 Evaluation
Figure 3: We generate the head models with randomly sampled identity codes and expression codes as condition. Each row corresponds to the same identity code, and each column corresponds to the same expression code.
@@ 4.2 Evaluation
Figure 4: We compared our initialization strategy with using the vertices of FLAME model. The left side shows the rendered image, and the right side shows the positions of the Gaussian points.
@@
We show the visualization results in Figure 4, with the Gaussian model rendering image on the left and the Gaussian positions displayed as point clouds on the right. Our initialization strategy using the guiding geometry model can ensure that all the Gaussian points fall evenly on the actual surface of the model, thereby ensuring reconstruction quality. When using the FLAME model for the initialization, a large number of points wander inside or outside the actual surface of the model, causing noise or redundancy and leading the model to lose some high-frequency information and making it difficult to fully converge. We also perform a quantitative evaluation of different initialization strategies on the rendered images, as shown in Table 1, which shows that our method leads to better rendering results.
@@
Table 1: Quantitative evaluation results of our initialization strategy and naive FLAME initialization strategy.
@@
Ablation on Representation and Super Resolution. We conduct the ablation study for the guiding mesh model, the Gaussian model and the super-resolution network (abbreviated as SR) as shown in the Fig. 5. The corresponding PSNR metrics are: Mesh (15.7), Mesh+SR (17.3), Gaussian (27.0), Gaussian+SR (29.3). Compared to mesh, utilizing 3D Gaussian as the representation brings significant improvements (+12), while the super resolution module adds some details, generating more realistic results.
Figure 5: The comparison of the different representations with super resolution.
4.3 Applications
@@
Image-based Fitting. In this section, we demonstrate the capability of our 3D Gaussian Parametric Model for single-image fitting using the fitting strategy detailed in Section 3.4. We compare our model with similar works: HeadNeRF [19], MoFaNeRF [64], and PanoHead [1]. In addition to evaluating the above methods on our evaluation dataset, we also conduct comparisons using cases from MEAD [46] dataset (the first two rows). The qualitative results are presented in Figure 6. Our model exhibits reconstruction accuracy while maintaining excellent 3D consistency and identity preservation. HeadNeRF’s fitting results often suffer from missing hair, and they remove the body and neck. MoFaNeRF, trained solely on the FaceScape dataset where all subjects wear hats, struggles to fit hair. As a GAN-based model, PanoHead can achieve highly accurate reproductions from the input view. However, due to overfitting, the results from side views reveal poor 3D consistency and identity preservation.
In addition to qualitative evaluations, we also conducted quantitative evaluations on 60 images using three metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), and Face Distance (FD). Here, we provide a brief explanation of the Face Distance (FD). To compute the FD metric, we utilized a face recognition tool  [^1] https://github.com/ageitgey/face_recognition to encode two images containing faces into 128-dimensional vectors. Subsequently, we calculated the distance between these two vectors to reflect the similarity of the two faces. In our experiments, FD serves as an indicator of identity consistency. The results are shown in Table 2. Our model demonstrates optimal performance in both fitting accuracy and identity consistency.
@@
Table 2: Quantitative evaluation results on the task of single image fitting. We compare our method with other 3 SOTA methods: HeadNeRF [19], MoFaNeRF [64], PanoHead [1].
@@
Expression Editing. Our 3D Gaussian Parametric Head Model possesses the capability for expression editing. Upon completing the fitting process on a portrait image, we can animate the model by applying different expression codes. The detailed pipeline is outlined in Section 3.4. An example is illustrated in Figure 7. Our model can generate images depicting the corresponding expressions of the input subject based on a reference expression (as seen in the lower left corner of each image in the figure). It performs admirably even with exaggerated expressions, producing natural and realistic results.
@@ 5 Discussion
Conclusion. In this paper, we propose the 3D Gaussian Parametric Head Model, a novel framework for parametric head model. This model leverages the power of 3D Gaussians, enabling realistic rendering quality and real-time speed. Our well-designed training strategy ensured stable convergence while enabling the model to learn appearance details and expressions. Besides, our model allows for creating detailed, high-quality face avatars from a single input image, and also enables editing for expressions and identity. We believe our model represents a significant advancement in the field of parametric head model.
Acknowledgements
The work is supported by the National Science Foundation of China (NSFC) under Grant Number 62125107 and the Postdoctoral Fellowship Program of China Postdoctoral Science Foundation under Grant Number GZC20231304.

### v2

@@ 4.1 Datasets
NeRSemble dataset contains over 260 different identities, and collects 72fps multi-view videos from 16 synchronized cameras for each identity. The combined video frames for each identity range from approximately 6000 to 11000 frames. For each identity video, we selected about 800 frames from all 16 views as training data.
NPHM dataset contains 5200 3D human head scans. These scans come from 255 different identities, each with about 20 different expressions. Since our method utilizes 2D images as training supervision, we render each scan from 80 different views to generate synthetic image data and record the camera parameters and the masks.
FaceVerse dataset is an East Asian human head scan dataset. It contains 2310 scans from 110 different identities, and each identity contains 21 expressions. We selected 1620 scan data of 80 identities for training. Similarly, for each scan, we render multi-view synthetic image data from 80 different views and record the camera parameters and the masks.
VFHQ dataset is a large-scale high-quality monocular video dataset, containing interviews and speeches of various people, with a total of 15,000+ videos, each with hundreds of frames. We removed some clips with poor quality and poor background segmentation, and selected 6,000 videos for training our model.
4.2 Evaluation for Gaussian Parametric Model.
Figure 5: We generate the head models with randomly sampled identity codes and expression codes as conditions. Each row corresponds to the same identity code, and each column corresponds to the same expression code.
@@
Figure 6: We compare our initialization strategy with using the vertices of FLAME model. The left side shows the rendered image, and the right side shows the positions of the Gaussian points.
@@
We show the visualization results in Fig. 6, with the Gaussian model rendering image on the left and the Gaussian positions displayed as point clouds on the right. Our initialization strategy using the guiding geometry model can ensure that all the Gaussian points fall evenly on the actual surface of the model, thereby ensuring reconstruction quality. When using the FLAME model for the initialization, a large number of points wander inside or outside the actual surface of the model, causing noise or redundancy and leading the model to lose some high-frequency information and making it difficult to fully converge. We also perform a quantitative evaluation of different initialization strategies on the rendered images, as shown in Table I, which shows that our method leads to better rendering results.
@@
TABLE I: Quantitative evaluation results of our initialization strategy and naive FLAME initialization strategy.
@@
Ablation on Representation and Super Resolution. We conduct the ablation study for the guiding mesh model, the Gaussian model, and the super-resolution network (abbreviated as SR) as shown in Fig. 7. The corresponding PSNR metrics are: Mesh (15.7), Mesh+SR (17.3), Gaussian (27.0), Gaussian+SR (29.3). Compared to mesh, utilizing 3D Gaussian as the representation brings significant improvements (+12), while the super-resolution module adds some details, generating more realistic results.
Figure 7: The comparison of the different representations with super-resolution.
4.3 Applications: Head Avatar from a Monocular Video.
Self Reenactment. We conduct qualitative and quantitative rendering quality comparisons between our method and five other SOTA monocular head avatar reconstruction methods on self reenactment task. Among them, AvatarMAV [3], NeRFBlendShape [4] and INSTA [5] utilize Voxel-based representation for NeRF head avatar training acceleration. FlashAvatar [29] and SplattingAvatar [28] introduce 3D Gaussians and bind them to a FLAME template to model the head avatars. In the experiment, we input 1-minute videos as training data and use additional 20-second videos as evaluation data. And we train their models according to the time claimed by each method. The training time of our method is 5 minutes. The qualitative results are shown in the Fig. 8. Our method is significantly better than other methods in terms of rendering quality, expression transfer accuracy and robustness. Table. II shows the quantitative evaluation results. We evaluate these methods on three metrics: Peak Signal-to-Noise Ratio (PSNR), Structure Similarity Index (SSIM) and Learned Perceptual Image Patch Similarity (LPIPS). Our method achieves the best results on all the metrics and significantly outperforms other SOTA methods on the LPIPS metric.
Figure 8: Qualitative comparison of our method and 5 other state-of-the-art methods on self reenactment task. From left to right: AvatarMAV [3], NeRFBlendShape [4], INSTA [5], FlashAvatar [29], SplattingAvatar [28] and Ours.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
AvatarMAV 27.7 0.92 0.081
NeRFBlendShape 27.9 0.92 0.085
INSTA 27.5 0.92 0.076
FlashAvatar 28.8 0.93 0.051
SplattingAvatar 28.4 0.93 0.065
Ours (wo FT motion) 28.1 0.93 0.046
Ours (wo $\boldsymbol{f_{dyn}}$) 28.2 0.93 0.044
Ours 28.9 0.94 0.041
TABLE II: Quantitative evaluation results on the task of self reenactment. We compare our method with other 5 SOTA methods: AvatarMAV [3], NeRFBlendShape [4], INSTA [5], FlashAvatar [29] and SplattingAvatar [28]. And we also include two ablation baselines: Ours (wo $\boldsymbol{f_{dyn}}$) in which we remove the dynamic generator $\boldsymbol{f_{dyn}}$, and Ours (wo FT motion) in which we only optimize the 3D Gaussians but the motion related networks.
Cross-identity Reenactment. We also compare our method with these state-of-the-art methods on the cross-identity reenactment task. The qualitative results are shown in Fig. 9. Other methods suffer from shape and expression coupling because they use a 3DMM model to control their avatars or use the 3DMM expression coefficient as a condition. As a result, the quality of the results is affected by the different shapes when applying cross-identity reenactment. Our method directly inputs the image into a well-decoupled expression encoder which is trained on large-scale datasets to extract latent expressions. Therefore, our method achieves better performance on cross-identity reenactment tasks.
Figure 9: Qualitative comparison of our method and 5 other state-of-the-art methods on cross-identity reenactment task. From left to right: AvatarMAV [3], NeRFBlendShape [4], INSTA [5], FlashAvatar [29], SplattingAvatar [28] and Ours.
Figure 10: We qualitatively compare our method with 5 other state-of-the-art methods on self reenactment tasks in the 100-shot, 10-shot, and 3-shot cases from top to bottom. The methods are AvatarMAV [3], NeRFBlendShape [4], INSTA [5], FlashAvatar [29], SplattingAvatar [28] and Ours from left to right.
Ablation on Few-shot Input. Next, we conduct experiments under the setting of few-shot input. We limit the number of input images to 100, 10, and 3 frames, and compare our method with the state-of-the-art methods mentioned above. The qualitative results are shown in Fig. 10. While other methods suffer from blurring, artifacts and significant quality degradation as the number of input images decreases, our method can achieve robust and high-quality avatar reconstruction. Even when inputting only one single image, our method can still guarantee robust and high-quality results as shown in Fig. 11.
Figure 11: The head avatar reconstruction result of our method when only one single image is used as input.
Ablation on Utilizing Synthesized Expression Condition Images.
As explained in Sec. 3.1 and Sec. 3.4, for the images input to the facial expression encoder, we utilize LivePortrait [30] to generate additional expression condition images. This strategy forces the encoder to learn only expression information from the input images, thereby achieving expression and appearance decoupling. We also construct an ablation baseline, in which we train the GPHM model using only the groundtruth images as the expression condition and the results are shown in Fig. 12. Without utilizing the additional expression condition images, the expression encoder leaks appearance-related information to the motion networks, causing the reconstructed head avatar to still be heavily affected by the actor’s appearance during cross-identity reenactment.
Figure 12: We use synthesized images via LivePortrait to train our expression encoder for expression and appearance decoupling.
@@
4.4 Applications: Image Fitting.
In this section, we demonstrate the capability of our 3D Gaussian Parametric Model for single-image fitting using the fitting strategy detailed in Section 3.5. We compare our model with similar works: HeadNeRF [21], MoFaNeRF [20], and PanoHead [45]. In addition to evaluating the above methods on our evaluation dataset, we also conduct comparisons using cases from MEAD [81] dataset (the first two rows). The qualitative results are presented in Fig. 13. Our model exhibits reconstruction accuracy while maintaining excellent 3D consistency and identity preservation. HeadNeRF’s fitting results often suffer from missing hair, and they remove the body and neck. MoFaNeRF, trained solely on the FaceScape dataset where all subjects wear hats, struggles to fit hair. As a GAN-based model, PanoHead can achieve highly accurate reproductions from the input view. However, due to overfitting, the results from side views reveal poor 3D consistency and identity preservation.
In addition to qualitative evaluations, we also conducted quantitative evaluations on 60 images using three metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), and Face Distance (FD). Here, we provide a brief explanation of the Face Distance (FD). To compute the FD metric, we utilized a face recognition tool  [^1] https://github.com/ageitgey/face_recognition to encode two images containing faces into 128-dimensional vectors. Subsequently, we calculated the distance between these two vectors to reflect the similarity of the two faces. In our experiments, FD serves as an indicator of identity consistency. The results are shown in Table III. Our model demonstrates optimal performance in both fitting accuracy and identity consistency.
@@
TABLE III: Quantitative evaluation results on the task of single image fitting. We compare our method with other 3 SOTA methods: HeadNeRF [21], MoFaNeRF [20], PanoHead [45].
@@
Our 3D Gaussian Parametric Head Model possesses the capability for expression editing. Upon completing the fitting process on a portrait image, we can animate the model by applying different expression codes. An example is illustrated in Figure 14. Our model can generate images depicting the corresponding expressions of the input subject based on a reference expression (as seen in the lower left corner of each image in the figure). It performs admirably even with exaggerated expressions, producing natural and realistic results.
@@ 5 Discussion
Conclusion. In this paper, we propose the 3D Gaussian Parametric Head Model, a novel framework for parametric head model. This model leverages the power of 3D Gaussians, enabling realistic rendering quality and real-time speed. Our well-designed training strategy ensures stable convergence while enabling the model to learn appearance details and expressions. Besides, our model allows for creating detailed, high-quality face avatars from a single input image, and also enables editing for expressions and identity. We believe our model represents a significant advancement in the field of parametric head model.
Acknowledgment The work is supported by the National Natural Science Foundation of China (NSFC) under Grant Number 62125107, 62402274 and the Postdoctoral Fellowship Program of China Postdoctoral Science Foundation under Grant Number GZC20231304.